2x - 7 + 4x = 4x - 10
Answer by nyc_function(2741)

2x - 7 + 4x = 4x - 10
6x - 7 = 4x - 10
6x - 4x = -10 + 7
2x = -3
x = -3/2
Did you follow?
By the way, this is a linear equation not linear algebra. Linear algebra is advanced algebra required only for math and engineering majors after completing calculus 1, 2 and 3.